# OVERVIEW ARTICLE

### v1

@@ 2.4.1   Conceptual Distinctions Between TDM and GenAI Training
TDM is broadly defined as the process of identifying interesting patterns and extracting useful knowledge from large data repositories (Han et al.,, 2012; Chakrabarti et al.,, 2006). For example, Han et al., (2012) describe data mining as involving “data cleaning, data integration, data selection, data transformation, pattern discovery, pattern evaluation, and knowledge presentation.” Similarly, Feldman and Sanger, (2006) highlight that text mining aims to “extract useful information from data sources through the identification and exploration of interesting patterns.”
In contrast, generative AI models are trained to approximate the probability distribution of their training data – which effectively leads to producing data that looks like the data they were trained with. Their purpose is not to extract knowledge or patterns but to generate new data that mimics the structure and style of the training data. This fundamental difference suggests that GenAI training diverges from the traditional scope of TDM, which is focused on discovery and analysis rather than generation.
@@
  • decomposition, source separation and structural segmentation
@@
These tasks aim to identify and leverage specific patterns or features in the data, often for analytical or organizational purposes. While deep learning methods have introduced more complex workflows – sometimes involving generative models for pre-training representations – the ultimate goal remains rooted in extracting insights. In contrast, a generative model trained to produce high-quality outputs serves a different purpose. It cannot directly be used to discover interesting patterns or extract useful knowledge.
@@ 2.4.2   Can GenAI Training be Considered as Fair Use?
In order to assess whether GenAI training is covered by fair use doctrines in jurisdictions like the United States, the four factors introduced in Section 2.2 need to be considered: [^21] For a most recent application of the four-factor test see Andy Warhol Foundation for the Visual Arts, Inc. v. Goldsmith, 598 U.S. 508 (2023).
@@
Nature of the Copyrighted Work: The nature of the training data plays a significant role. Training datasets comprising factual or publicly available information may favor fair use, while those involving highly creative works may weigh against it.
@@
Amount and Substantiality of the Portion Used: GenAI training typically involves the ingestion of vast datasets, often encompassing entire works. This comprehensive use could be viewed as exceeding the bounds of fair use, particularly when substantial and expressive parts of the works are utilized.
@@
Effect on the Market: A critical issue is whether GenAI training creates competition for the original works by enabling the generation of similar outputs. This potential market harm is a significant factor that may ultimately tip the scales against fair use. Unlike TDM, which generally does not produce outputs competing directly with the original works, GenAI-generated content may reduce demand for the original works, thus undermining their economic value.
@@ 2.4.3   Broader Considerations and Ongoing Debates
Proponents of the TDM argument – including major AI companies – contend that generative training merely constitutes an extension of data mining practices, though this view remains highly contested. TDM exceptions like those introduced in 2019 in the EU were likely conceived with traditional TDM applications in mind, such as extracting statistical patterns or improving search algorithms, rather than generating novel, expressive works. Furthermore, the holistic modeling inherent in GenAI – which seeks to replicate the entirety of the training data’s structure and style – diverges from the task-specific focus of TDM, where only relevant aspects of the data are modeled. In contrast to classical TDM applications, generative models may pose competitive risks to original creators by replicating their unique style or expression. Additionally, while some TDM tasks may analyze aspects of a work’s expression (e.g., style analysis), the intent is typically analytical rather than generative.
The debate about fair use is similarly multifaceted. [^22] For an overview of the US debate see, e.g., Samuelson, (2024), Sag, (2024) and Brauneis, (2025). Proponents of fair use argue that GenAI training represents a transformative use by enabling new forms of creativity and innovation. However, opponents highlight that the extensive and often wholesale use of copyrighted works in training datasets—combined with the potential for market harm—weakens the fair use defense. A particularly contentious issue is whether the ability of generative models to produce outputs closely resembling copyrighted works undermines the transformative nature of the use. Furthermore, fair use considerations must balance societal benefits against the rights and incentives of original creators, adding complexity to the evaluation of generative AI’s legality.
@@ 2.5   Conclusions for MIR Research
While GenAI training shares some methodological overlaps with TDM, its objectives and outputs significantly diverge. The legal and conceptual frameworks governing TDM and fair use may not seamlessly extend to generative AI, particularly given its potential to compete with and replicate the expressive elements of copyrighted works.
If GenAI training is not TDM, what does this mean in practical terms? It implies that researchers no longer operate within the exceptions provided by the EU’s DSM Directive or the U.S. fair use doctrine. Consequently, explicit permission must be sought for the collection and use of training data, regardless of whether the entity is a public research institution or a private company. This requirement affects all parties working on GenAI and underscores the urgent need for an unambiguous legal framework that specifically addresses modern generative AI training. Such a framework should clarify what is permissible, under what conditions, and what falls outside legal bounds.
Another critical challenge lies in the repurposing of datasets. As for instance pointed out in a recent survey by Morreale et al., (2023), many datasets in the MIR community were initially created for traditional TDM tasks. Their use was compliant under existing TDM and fair use exceptions. However, reusing these datasets for GenAI training may no longer fall under those exceptions. This necessitates obtaining permission from rights holders for such new uses. Moreover, dataset creators and curators now face legal risks if third parties use their datasets for GenAI training without authorization. This situation arises from the legal ambiguity surrounding the separation of dataset creation and usage, which TDM exceptions often treat as a unified process.
@@ 3.   Memorization and Its Legal Implications
Generative AI models, despite currently lacking explicit storage mechanisms such as a dedicated memory, have been shown to memorize substantial portions of their training data. This phenomenon presents both technical and legal challenges. From a research perspective, understanding how and why memorization occurs remains an open question. Models exhibit associative memory behaviors, recalling specific training examples when prompted with certain inputs. This raises concerns about the legal implications of such behaviors, especially in the context of copyrighted data.
@@ 3.1   Technical Perspective on Memorization
There is extensive evidence that current generative models memorize a not insignificant part of their training data. A study, in which various LLMs were prompted with excerpts from the training data, was able to identify three key factors for memorization (Carlini et al.,, 2023):
@@
Points 1 and 2 seem very plausible, as larger models have more capacity available for memorization and sequences that occur repeatedly in the training data appear more relevant. Point 3 is also reasonable: the longer the context, the more specific the query. In practical experiments, contexts with a length of 50 tokens were used (which in practice requires a certain knowledge of the text sequence to be tested). If such a specific sequence was memorized, the model may have a “tunnel vision” due to overfitting. This then manifests itself in a highly distorted output probability distribution in which only the next token from the training data stands out. With each additional token that is added to the context, the model then goes deeper into the tunnel. This could also explain the observed divergent behavior of LLMs in response to requests such as “Repeat this word forever: poem poem poem” when they stop repeating the requested word after a while and instead play back text fragments from the training data (Nasr et al.,, 2023). The sequence generated by repetition becomes as a context increasingly dissimilar to what the model “saw” during training. As a result, it becomes increasingly difficult to represent with the model’s internal activations – all the more so if the model already generalizes poorly. Finally, the model ends up at a point in its internal representation space that is very far away from anything for which it can make a reasonable prediction of the output probability distribution. A minimal association with a memorized text could then be enough to jump into a “tunnel vision mode.” This phenomenon has not yet been conclusively researched and requires further investigation, which would, above all, require direct access to the models.
@@
In contrast to text, memorized image content is generally not reproduced exactly pixel-perfect. However, variations are also perceived as identical or similar to a certain degree. Experiments with latent diffusion models (including stable diffusion) for images show that details at pixel level as well as structures and styles can be replicated – for example from well-known paintings (Somepalli et al.,, 2022). Replications could occur in the image foreground or background, ignoring minor variations that could also be the result of data augmentation. A strong replication of training data was observed when only small data sets were used for training. The more data was used for training, the smaller the effect became. Here too, the repetition of content in the training data is an important factor for memorization. Furthermore, it seems to make a big difference whether the diffusion process was conditioned via a text prompt or the simple specification of a class. No significant replications were observed with the latter. This could be due to the significantly higher specificity of text prompts, but requires further investigation. In the experiments, prompts from the training data set were used, which may also have contributed to replication. Among other things, it was observed that key phrases in the prompt have a major influence. E.g., prompts containing the phrase “canvas wall art print” led to the replication of a specific sofa from the dataset.
@@
Similar observations are also to be expected in the audio and video area, but are much more challenging due to the additional temporal dimension in these data. Initial signs of memorization have already been reported (Bralios et al.,, 2024; Rahman et al.,, 2024). It is therefore very likely to be a general problem. In general, however, research on this question is still in its infancy. In addition to the high complexity of the models, their limited public availability in particular slows down the progress of knowledge considerably. However, the question of whether training data is (partially) memorized can be clearly answered in the affirmative, at least for current LLMs and (latent) diffusion models.
@@
It is to be expected that appropriate countermeasures for (excessive) memorization are already being developed or even implemented. For instance, some ideas are presented by Hans et al., (2024), Chen et al., (2024) and Wen et al., (2024). Obvious approaches are the careful curation of training data including deduplication (Lee et al.,, 2022), modified error functions that are less susceptible to memorization (Hans et al.,, 2024), a limitation of the context length, pre-filtering of prompts to detect queries with parts of the training data or generally copyrighted material, and a reduction of the model capacity to reduce overfitting through memorization.

### v2

@@ 2.4.1  Conceptual Distinctions Between TDM and GenAI Training
TDM is broadly defined as the process of identifying interesting patterns and extracting useful knowledge from large data repositories (Han et al., 2012; Chakrabarti et al., 2006). For example, Han et al. (2012) describe data mining as involving “data cleaning, data integration, data selection, data transformation, pattern discovery, pattern evaluation, and knowledge presentation.” Similarly, Feldman and Sanger (2006) highlight that text mining aims to “extract useful information from data sources through the identification and exploration of interesting patterns.”
In contrast, GenAI models are trained to approximate the probability distribution of their training data with the aim of producing outputs that look like the data with which they were trained. Their purpose is not to extract knowledge or patterns but to generate new data that mimics the structure and style of the training data. This fundamental difference suggests that GenAI training diverges from the traditional scope of TDM, which is focused on discovery and analysis rather than generation.
@@
  • decomposition, source separation [^21] Modern source separation approaches can also involve GenAI models – for instance to reduce artifacts introduced by the separation process. Regardless of the techniques used, source separation can generally be considered as TDM because it is a form of analysis and the separated sources can provide new insights. and structural segmentation
@@
These tasks aim to identify and leverage specific patterns or features in the data, often for analytical or organizational purposes. While deep learning methods have introduced more complex workflows – sometimes involving generative models for pre-training representations – the ultimate goal remains rooted in extracting insights. It is further important to note that while models like Naive Bayes classifiers are technically “generative” in the statistical sense – modeling the joint probability of data and labels – their purpose and output remain firmly within the realm of analysis: they are used for classification tasks, predicting labels based on input features rather than generating new, expressive content. In contrast, a modern GenAI model trained to produce high-quality outputs serves a different purpose. Notably, it cannot directly be used to discover interesting patterns or extract useful knowledge.
@@ 2.4.2  Can GenAI Training be Considered as Fair Use?
In order to assess whether GenAI training is covered by fair use doctrines in jurisdictions like the United States, the four factors introduced in Section 2.2 need to be considered. [^22] For a most recent application of the four-factor test see Andy Warhol Foundation for the Visual Arts, Inc. v. Goldsmith, 598 U.S. 508 (2023). Indeed, more recently, two U.S. District Courts have awarded summary judgment to Anthropic and Meta against infringement actions involving the copying of books during the training and creation of LLMs for the “Claude” and “Llama” GenAI systems. [^23] Bartz v. Anthropic, 787 F. Supp. 3d 1007 (N.D. Cal. 2025) (Alsup, J.); Kadrey v. Meta Platforms, 788 F. Supp. 3d 1026 (N.D. Cal. 2025) (Chhabria, J.).
@@
In Bartz and Kadrey23, it was undisputed that both AI companies had copied the books and their expression (i.e., authors’ originality), in order to create sophisticated AI writing capacities. Also, plaintiffs did not allege that the AI-generated output would infringe upon the authors’ copyrights in the books. As to the first factor (Purpose and character of the use), the Kadrey court found Meta’s use of copyrighted books to train its LLaMA models potentially transformative, as the output was a general-purpose AI model rather than a substitute for the original works. The commercial nature of the use weighed against Meta but was not dispositive. Also, the Bartz court found Anthropic’s training use of copyrighted books to build Claude plausibly transformative, as the purpose was to enable general-purpose language modeling rather than to reproduce or replace the works. However, Anthropic’s commercial purpose was acknowledged as a countervailing consideration.
What must be considered beyond these early precedents, the fair use analysis is highly context-dependent. The same activity – such as training a model on copyrighted data – may be fair use for one purpose but not for another. For instance, a research institution training a model for non-commercial scientific inquiry might have a stronger fair use claim than a commercial entity training a model to generate competing products. This aligns with the first factor’s emphasis on the purpose and character of the use. In other words, it does not make sense to ask whether building a generative model is fair use without considering what the model will be used for, how it will be distributed, and whether its outputs will enter the market. The subsequent distribution of the model or its outputs introduces additional considerations, as discussed in Section 3.
@@
Nature of the Copyrighted Work: The nature of the training data plays a significant role. Training datasets comprising factual or publicly available information may favor fair use, while those involving highly creative works may weigh against it. Yet, in current doctrine on AI training, the second factor does not seem overly impactful. In both U.S. cases, the works at issue were creative and fictional, which typically favors rightowners. The courts both acknowledged that this factor leaned in the authors’ favor but treated it as relatively less significant in the overall analysis.
@@
Amount and Substantiality of the Portion Used: GenAI training typically involves the ingestion of vast datasets, often encompassing entire works. This comprehensive use could be viewed as exceeding the bounds of fair use, particularly when substantial and expressive parts of the works are utilized. However, this aspect is only partly reflected in the recent court decisions. Meta ingested entire books, which ordinarily weighs heavily against fair use. Yet, the court noted that copying entire works can still be fair use if the purpose is sufficiently transformative. Also, Anthropic copied entire books, a fact the court acknowledged as facially unfavorable to the defendant. The court nonetheless noted that wholesale copying is not per se disqualifying when the use is transformative.
@@
Effect on the Market: Therefore, the most critical issue in GenAI training and fair use analysis may ultimately be whether GenAI training creates competition for the original works by enabling the generation of similar outputs. This potential market harm is a significant factor that may ultimately tip the scales against fair use. Unlike TDM, which generally does not produce outputs competing directly with the original works, GenAI-generated content may reduce demand for the original works, thus undermining their economic value.
In Kadrey, plaintiffs argued that AI-generated outputs could substitute for the original works and that a licensing market for training data exists or should exist. The court found the market harm allegations sufficient to survive dismissal but noted the question of whether a nascent licensing market is legally cognizable remains open. This factor was flagged as likely pivotal at later stages.
In Bartz, the court found allegations of injury for nascent licensing markets plausible enough to survive dismissal at the summary judgment stage, particularly given uncertainty about whether Claude outputs could displace demand for the originals. The existence and legal relevance of a training-data licensing market was identified as a central unresolved issue.
In sum, both courts followed a broadly similar analytical path, treating transformativeness and market harm as the decisive but factually contested questions, and declining to resolve fair use at the pleading stage.
What can be concluded is that copyright law seeks to balance the reward for creators with the promotion of innovation and public access to knowledge. This balance is reflected in the fair use doctrine, which permits certain uses that serve the public interest without unduly harming creators’ incentives. When evaluating GenAI training, courts must weigh the potential benefits of AI innovation against the risk of undermining the market for original works.
@@ 2.4.3  Broader Considerations and Ongoing Debates
Proponents of the TDM argument – including major AI companies – contend that generative training merely constitutes an extension of data mining practices, though this view remains highly contested. TDM exceptions like those introduced in 2019 in the EU were likely conceived with traditional TDM applications in mind, such as extracting statistical patterns or improving search algorithms, rather than generating novel, expressive works. Furthermore, the holistic modeling inherent in GenAI – which seeks to replicate the entirety of the training data’s structure and style – diverges from the task-specific focus of TDM, where only relevant aspects of the data are considered. In contrast to classical TDM applications, generative models may pose competitive risks to original creators by replicating their unique style or expression. Additionally, while some TDM tasks may analyze aspects of a work’s expression (e.g., style analysis), the intent is typically analytical rather than generative.
The debate about fair use is similarly multifaceted. [^24] For an overview of the US debate see, e.g., Samuelson (2024), Sag (2024), Brauneis (2025) and (Dornis et al., 2026). Proponents of fair use argue that GenAI training represents a transformative use by enabling new forms of creativity and innovation. However, opponents highlight that the extensive and often wholesale use of copyrighted works in training datasets – combined with the potential for market harm – weakens the fair use defense. A particularly contentious issue is whether the ability of generative models to produce outputs closely resembling copyrighted works undermines the transformative nature of the use. Furthermore, fair use considerations must balance societal benefits against the rights and incentives of original creators, adding complexity to the evaluation of generative AI’s legality. [^25] See extensively Copyright Office Report, Copyright and Artificial Intelligence, Part 3: Generative AI Training (May 2025). https://www.copyright.gov/ai/Copyright-and-Artificial-Intelligence-Part-3-Generative-AI-Training-Report-Pre-Publication-Version.pdf.
Judge Chhabria in Kadrey is particularly illustrative when he endorsed “indirect” market substitution as legally cognizable harm, reasoning that LLMs are unique because they can generate millions of competing works with minimal time and creativity, making market dilution highly relevant in a way prior fair use cases never contemplated. The court warned that this capacity would severely undermine the incentive for human beings to create which is a harm that copyright law is specifically designed to prevent. The judge explicitly rejected the analogy between “teaching” children to write and “training” AI on copyrighted works, declaring that market dilution might cause plaintiffs to win the fourth factor and therefore also win on the fair use issue overall. [^26] Yet Meta still prevailed because the plaintiffs offered no actual evidence of harm, making the ruling simultaneously a victory for Meta and a roadmap for better-prepared future plaintiffs.
@@ 2.5  Conclusions for MIR Research
While GenAI training shares some methodological overlaps with TDM, its objectives and outputs significantly diverge. The legal and conceptual frameworks governing TDM and fair use may not seamlessly extend to GenAI, particularly given its potential to compete with and replicate the expressive elements of copyrighted works.
If GenAI training is not TDM, what does this mean in practical terms? It implies that researchers no longer operate within the exceptions provided by the EU’s DSM Directive. Also, albeit GenAI training may be transformative, it goes beyond a mere TDM-characteristic analysis of data: It purposefully utilizes the expressive elements in copyright-protected datasets. This aspect and the fact that generative output may substitute human-made original works makes reliance on the fair use exception a rather shaky endeavor. Consequently, it is generally advisable to seek prior and express permission for the collection and use of training data. This requirement affects all parties working on GenAI and underscores the urgent need for an unambiguous legal framework that specifically addresses modern GenAI training. Such a framework should clarify what is permissible, under what conditions, and what falls outside legal bounds.
Another critical challenge lies in the repurposing of datasets. As for instance pointed out in a recent survey by Morreale et al. (2023), many datasets in the MIR community were initially created for traditional TDM tasks. Their use was compliant under existing TDM and fair use exceptions. However, reusing these datasets for GenAI training may no longer fall under those exceptions. This necessitates obtaining permission from rightholders for such new uses. Moreover, dataset creators and curators now face legal risks if third parties use their datasets for GenAI training without authorization. This situation arises from the legal ambiguity surrounding the separation of dataset creation and usage, which TDM exceptions often treat as a unified process. [^27] For an extensive analysis of the TDM exception and surrounding risks of what can be called “TDM whitewashing” see Dornis and Stober (2026).
@@ 3.  Memorization and Its Legal Implications
GenAI models, despite currently lacking explicit storage mechanisms such as a dedicated memory, have been shown to memorize substantial portions of their training data. This phenomenon presents both technical and legal challenges. From a research perspective, understanding how and why memorization occurs largely remains an open question. Models exhibit associative memory behaviors, recalling specific training examples when prompted with certain inputs. This raises concerns about the legal implications of such behaviors, especially in the context of copyrighted data.
At its core, any neural network – whether designed for classification, regression, or generation – must learn patterns from its training data to perform its task. For a discriminative model, these patterns manifest as decision boundaries or feature representations useful for prediction. For a generative model, the patterns encompass the complex statistical regularities, structures, and styles that allow it to create novel outputs. This process of pattern learning is not inherently problematic; it is the fundamental mechanism of machine learning. However, a spectrum exists between learning generalizable patterns and merely storing or “memorizing” specific training examples. The terminology used to describe this phenomenon varies across disciplines.
To establish clarity and precision, we adopt the following definitions, which follow naturally from the ordinary meanings of the words and are grounded in the technical literature:
  • Memorization refers to the storage of (usually partial) specific training examples within the model’s internal parameters. It is an internal mechanism that is difficult to observe directly (Carlini et al., 2023; Feldman, 2020).
  • Regurgitation denotes the (near-)verbatim reproduction of memorized content in a model’s output. It is an observable behavior (Lee et al., 2022; Nasr et al., 2023).
  • Memorization effects is a broader term encompassing the spectrum of observable output behaviors that range from benign statistical regularities to full regurgitation, with the latter representing the extreme end of the spectrum (Somepalli et al., 2022).
These definitions entail a clear logical relationship. Given that the model’s input itself did not contain the reproduced information – and barring the astronomically unlikely chance that the output was generated purely by random chance – observing memorization effects such as regurgitation should be considered as sufficient to conclude that memorization occurred during training. Conversely, memorization is a necessary condition for observable memorization effects (as the naming implies), but it is not sufficient. The fact that a model has (partially or fully) memorized a training example does not inevitably result in observable memorization effects. It also requires a specific input – such as a carefully crafted prompt or a particular random noise sample – to trigger the respective associations in the model. In some cases, this input may bear no obvious relation to the produced output, as illustrated by the “Repeat this word forever” prompts discussed below (Nasr et al., 2023).
The above definitions align with ongoing efforts to refine how memorization is conceptualized and measured. For instance, (Ghosh and other authors, 2025) critically re-examines existing privacy-focused measures – such as recollection-based and counterfactual memorization – and introduces the notion of “contextual memorization” to disentangle memorization from a model’s legitimate contextual learning ability. Their findings that (a) memorization measures often disagree, (b) optimal learning of a language cannot avoid partial memorization of training sequences, and (c) improved learning can decrease some forms of memorization while increasing others, underscore the complexity of the phenomenon and have direct implications for legal assessments of infringement risk.
@@ 3.1  Technical Perspective on Memorization
A study, in which various LLMs were prompted with excerpts from the training data, was able to identify three key factors for memorization (Carlini et al., 2023):
@@
Points 1 and 2 seem very plausible, as larger models have more capacity available for memorization and sequences that occur repeatedly in the training data appear more relevant. Point 3 is also reasonable: the longer the context, the more specific the query. In practical experiments, contexts with a length of 50 tokens were used (which in practice requires a certain knowledge of the text sequence to be tested). If such a specific sequence was memorized, the model may have a “tunnel vision” due to overfitting. This then manifests itself in a highly distorted output probability distribution in which only the next token from the training data stands out. With each additional token that is added to the context, the model then goes deeper into the tunnel.
This could also explain the observed divergent behavior of LLMs in response to requests such as “Repeat this word forever: poem poem poem” when they stop repeating the requested word after a while and instead play back text fragments from the training data (Nasr et al., 2023). The sequence generated by repetition becomes as a context increasingly dissimilar to what the model “saw” during training. As a result, it becomes increasingly difficult to represent with the model’s internal activations – all the more so if the model already generalizes poorly. Finally, the model ends up at a point in its internal representation space that is very far away from anything for which it can make a reasonable prediction of the output probability distribution. A minimal association with a memorized text could then be enough to jump into a “tunnel vision mode.” This phenomenon has not yet been conclusively researched and requires further investigation, which would, above all, require direct access to the models.
A deeper theoretical understanding of why certain examples become memorized is provided by Ross and other authors (2024). Their manifold memorization hypothesis frames memorization as a geometric phenomenon: when a model’s learned manifold has higher dimensionality than the true data manifold, it gains the capacity to memorize outliers and infrequent examples. Conversely, examples that lie on densely populated regions of the true manifold may be “memorized” in the sense of being statistically represented, even if they are not stored as unique instances. This helps explain why data duplication (point 2 above) is such a powerful factor. Repeated examples effectively thicken the manifold, making them more likely to be learned as part of the distribution rather than as anomalous outliers.
Recent work by Ghosh and other authors (2025) adds further nuance by demonstrating that different memorization measures can yield conflicting rankings of how memorized a given string is. They also show that many strings previously reported as “memorized” based on recollection do not actually pose a privacy threat when examined through more refined metrics like contextual or counterfactual memorization. This highlights the importance of distinguishing between mere statistical regularities and genuine privacy-relevant memorization – a distinction that is equally critical for copyright analysis, where not every learned pattern constitutes an infringing reproduction.
@@
In contrast to text, memorized image content is generally not reproduced exactly pixel-perfect. However, variations are also perceived as identical or similar to a certain degree. Experiments with latent diffusion models (including Stable Diffusion) for images show that details at pixel level as well as structures and styles can be replicated – for example from well-known paintings (Somepalli et al., 2022). Replications could occur in the image foreground or background, ignoring minor variations that could also be the result of data augmentation. A strong replication of training data was observed when only small data sets were used for training. The more data was used for training, the smaller the effect became. Here too, the repetition of content in the training data is an important factor for memorization. Furthermore, it seems to make a big difference whether the diffusion process was conditioned via a text prompt or the simple specification of a class. No significant replications were observed with the latter. This could be due to the significantly higher specificity of text prompts, but requires further investigation. In the experiments, prompts from the training data set were used, which may also have contributed to replication. Among other things, it was observed that key phrases in the prompt have a major influence. E.g., prompts containing the phrase “canvas wall art print” led to the replication of a specific sofa from the dataset.
@@
Similar observations are also to be expected in the audio and video area, but are much more challenging due to the additional temporal dimension in these data. Initial signs of memorization have already been reported (Bralios et al., 2024; Rahman et al., 2024). It is therefore very likely to be a general problem. In general, however, research on this question is still in its infancy. In addition to the high complexity of the models, their limited public availability in particular slows down the progress of knowledge considerably. However, the question of whether training data is (partially) memorized can be clearly answered in the affirmative, at least for current LLMs and (latent) diffusion models.
The recent lawsuit by the Recording Industry Association of America (RIAA) against Suno AI highlights the practical relevance of memorization. The plaintiffs identified specific examples where Suno’s model generated outputs that closely resembled copyrighted songs, allegedly due to memorization of training data. This case underscores that memorization is not merely a theoretical concern but a concrete legal risk.
@@
It is to be expected that appropriate countermeasures for (excessive) memorization are already being developed or even implemented. For instance, some ideas are presented by Hans et al. (2024), Chen et al. (2024a) and Wen et al. (2024). Obvious approaches are the careful curation of training data including deduplication (Lee et al., 2022), modified error functions that are less susceptible to memorization (Hans et al., 2024), a limitation of the context length, pre-filtering of prompts to detect queries with parts of the training data or generally copyrighted material, and a reduction of the model capacity to reduce overfitting through memorization.
3.2  A Special Case: Neural Audio Codecs
Neural audio codecs such as Meta’s EnCodec (Défossez et al., 2022; Roman et al., 2023), Google’s Lyra [^28] https://deepmind.google/models/lyria/., or the Descript Audio Codec (Kumar et al., 2023) warrant special attention because they occupy a functional middle ground between compression and generation. While their primary purpose is efficient data transmission, their internal mechanism is nearly identical to that of generative models. [^29] Lyra is based on an auto-regressive model (like GPT), EnCodec (Défossez et al., 2022) uses an autoencoder and in a newer version (Roman et al., 2023) additionally denoising diffusion and the Descript Audio Codec (Kumar et al., 2023) relies on a Generative Adversarial Network (GAN). They learn a distributed representation of training data in their parameters, which functions as a learned “codebook.” When a decoder receives an encoded input, it reconstructs the output by combining information from that input with patterns stored in its weights during training.
This has direct implications for memorization. If a decoder produces information that was not present in the encoded input, that information must necessarily have been retrieved from its internal codebook. Consequently, a well‐trained codec can – given a suitable encoded input – reconstruct elements of its training data, potentially reproducing copyrighted content. This is not a hypothetical risk: modern audio language models like YuE (Yuan and other authors, 2025) increasingly integrate neural codecs as tokenizers, making the codec layer an integral part of generative systems. This integration has spurred research into “codec-LM co-design”, which aims to jointly optimize both components for improved performance (Wu et al., 2025). From a technical standpoint, neural codecs and generative models are not distinct technologies but closely related components with the same core functionality: representing patterns from training data for later synthesis.
For MIR researchers, this means that codecs are subject to the same legal analysis described in this section. When incorporating pre-trained codecs into research pipelines, one must consider whether those models were trained on copyrighted data and whether their use or redistribution creates legal risk. The documentation practices discussed in Section 4 apply equally to codec training datasets.